OpenStudy (anonymous):

A car starts from rest and accelerates uniformly at 3.0 m/s2. A second car starts from rest 6.0 seconds later at the same point and accelerates uniformly at 5.0 m/s2. How long does it take the second car to overtake the first car?

OpenStudy (anonymous):

I'm using 1/2at^2 = 1/2a(t-6)^2 but getting confused somewhere in the math.

OpenStudy (shane_b):

I don't see an issue with what you're doing...you should end up with a value of 26.6 secs. Maybe if you show your calcs someone can chime in a figure out what's going on (logging off soon).

OpenStudy (anonymous):

3/2t^2 = 5/2(t^2-12t+36) 3/2t^2 = 5/2t^2-30t+90 t^2 = 5/3t^2-20t+60 Still working on the rest!

OpenStudy (anonymous):

 you can think an other way X1 = 1/2a1*t^2+vo*t + xo , vo = 0 , xo = 1/2a1*t1*2=24 X1 = 1/2a1*t^2 + 24 X2 = 1/2a2 * t^2 second car over take first car unless X1 = X2 1/2a1*t^2 + 24 = 1/2a2*t2